# LiDARCrafter: Dynamic 4D World Modeling from LiDAR Sequences

### v1

@@ 3.3 Scene2Seq: Autoregressive LiDAR Synthesis
A core innovation of LiDARCrafter is its ability to generate the LiDAR stream autoregressively. In RGB video, textures and lighting change constantly, whereas a LiDAR sweep sees a mostly static environment, with only the ego vehicle and annotated objects moving. We exploit this stability by warping previously observed points to create a strong prior, as shown in fig. 4. Concretely, we back-project the first range image $\mathbf{I}^{0}$ to a point cloud $\mathbf{P}^{0}$, then split it with the layout boxes into background $\mathbf{B}^{0}$ and foreground sets $\{\mathbf{F}^{0}_{i}\}_{i=1}^{M}$. In later frames, we warp $\mathbf{B}^{0}$ with the ego pose, and update each $\mathbf{F}^{0}_{i}$ by its own motion prior, providing a strong drift-free geometric prior for the diffusion model at every denoising step.
@@
Dynamic-Object Warp. For each object $i$, we first shift its box center by its own cumulative planar offsets $(\Delta x_{i}^{t},\Delta y_{i}^{t})$, giving the world-frame position $\mathbf{u}_{i}^{t}=[x_{i}+\Delta x_{i}^{t},\;y_{i}+\Delta y_{i}^{t},\;z_{i}]^{\top}$. Its heading change $\psi_{i}^{t}$ is obtained exactly as in eq. 7. To express the box in the current ego frame, we apply the inverse ego transform: first translate by $\mathbf{u}_{0}^{t}$ and then rotate by $\psi_{0}^{t}$. The same rigid transform maps the stored foreground points $\mathbf{F}_{i}^{0}$ to $\mathbf{F}_{i}^{t}$. These warped foreground object points, combined with the updated background points, supply a strong geometric prior for the later timestep.
@@ 4 Experiments
Table 1: Evaluations of scene-level fidelity for LiDAR generation on the nuScenes dataset. MMD values are reported in $10^{-4}$ and JSD in $10^{-2}$. Lower is better for all metrics ($\downarrow$).
@@
Table 5: Comparison of object generation consistency using CFCA ($\uparrow$) and CFSC ($\uparrow$). CFCA measures classification accuracy on generated points using a PointMLP trained on real data. CFSC assesses geometric consistency by regressing boxes from generated points and computing IoU; the number indicates the point count within each box.
@@
Table 6: Comparison of temporal consistency in 4D LiDAR generation. TTCE measures transformation error from point cloud registration; CTC computes Chamfer Distance between adjacent frames. Numbers indicate frame intervals.
@@
Table 7: Ablation on foreground conditioning methods for LiDAR generation. 2D masks are projected from 3D boxes.
@@
4.1 Experimental Settings
We benchmark LiDARCrafter on the widely used nuScenes dataset (Caesar et al. 2020), whose 32-beam LiDAR sweeps, 3D boxes, instance IDs, and HD maps provide rich conditioning cues. Evaluation combines classic static scores (FRD, FPD, JSD, MMD) with our object-, layout-, and motion-centric metrics from section 3.4. Training uses six NVIDIA A40 GPUs, the three-branch layout diffuser runs $1M$ steps at batch $64$, and the range-image model ($32\times 1024$) runs $500K$ steps at batch $32$. All diffusers adopt $1024$ denoise steps for training and $256$ for sampling. More experimental settings appear in the appendix.
4.2 Scene-Level LiDAR Generation
Whole-Scene Fidelity. Table 1 compares LiDARCrafter against recent LiDAR generation baselines in terms of whole-scene fidelity. Our method outperforms all existing methods, achieving the lowest FRD ($194.37$, a $20\%$improvement over R2DM) and FPD ($8.64$), indicating more accurate reconstruction of both 2D range structures and 3D point distributions. As shown in Fig. 5, our method produces scans that closely resemble the ground truth, with well-preserved foreground structures, while other methods suffer from background noise or blurring.
Foreground Object Accuracy. To evaluate the fidelity of synthesized foregrounds, we apply a pre-trained VoxelRCNN detector (Deng et al. 2021) to the generated scenes. We report Foreground Detection Confidence (FDC), which reflects the detector’s confidence (cf. Table 2), and Conditioned Detection Accuracy (CDA), which measures the average precision (AP) of detected boxes (cf. Table 3). LiDARCrafter achieves the highest FDC scores across most categories, indicating superior object fidelity, and also attains the best detection AP, demonstrating stronger alignment between synthesized structures and the given conditions.
@@ 4.3 Object-Level LiDAR Generation
Semantic and Geometric Consistency. To further evaluate object quality under conditioning, we introduce two metrics: CFCA for semantic fidelity and CFSC for geometric consistency (Table 5). For semantic fidelity, we apply a PointMLP (Ma et al. 2022) classifier trained on real data to classify generated instances, yielding a CFCA score of $73.48\%$ for LiDARCrafter, indicating strong alignment with real-world categories. For geometric consistency, we use a conditional variational autoencoder to regress bounding boxes from generated point clouds, and compute the mean IoU with ground truth. LiDARCrafter achieves the highest IoU across all point count settings, demonstrating superior adherence to geometric constraints.
@@ 4.4 Autoregressive 4D LiDAR Generation
Temporal Consistency. We evaluate temporal consistency in 4D LiDAR generation in Table 6. TTCE measures the error between the predicted and ground-truth transformation matrices obtained via point cloud registration, while CTC computes the Chamfer Distance between consecutive frames. Our approach achieves the lowest TTCE scores across both frame intervals and maintains competitive CTC performance at all intervals, demonstrating strong temporal coherence. Qualitative comparisons in Fig. 6 further show that LiDARCrafter produces sequences with consistent structure and fine geometric detail, whereas other methods often suffer from degraded fidelity over time.
@@ 4.6 Applications
Leveraging its object-centric generation capability, LiDARCrafter enables the synthesis of rare and diverse corner-case scenarios, which are valuable for data augmentation and for evaluating the robustness of downstream algorithms. As illustrated in Fig. 7, our method can generate challenging situations such as a vehicle cutting into the driving lane, a pedestrian emerging in a bus blind spot, a collision between two adjacent cars, and an overtaking maneuver. These results demonstrate that our method not only produces the desired scenarios on demand, but also maintains strong temporal coherence throughout the sequence. Additional qualitative examples are provided in the appendix.
@@ 5 Conclusion
$\printcontents[appendices]$l1
Appendix A Additional Implementation Details
@@
  • Notation Summary (Sec. A.1): We summarize the notations used throughout the main manuscript for clarity.
  • Scene Graph Construction (Sec. A.2): We explain how scene graphs are built from dataset annotations, which serve as training data for LiDARCrafter.
  • Training Pipeline (Sec. A.3): We detail the data preprocessing steps, loss functions, and hyperparameter settings used for both 4D layout and point cloud generation.
  • Evaluation Suite (Sec. A.4): We clarify the motivation behind each proposed metric and specify the exact computation procedures.
  • Baseline Models (Sec. A.5): We briefly describe the baseline methods used for comparisons.
@@
A.1 Summary of Notatations
For better readability, the notations used in this work have been summarized in Tab. 9.
A.2 Scene Graph Construction from Annotations
To train LiDARCrafter, we construct an ego-centric scene graph for each LiDAR frame using annotations from the nuScenes dataset. Each scene graph consists of nodes representing the ego vehicle and nearby foreground objects, along with directed edges that capture their spatial and semantic relationships. The construction pipeline is summarized in Alg. 1, and detailed as follows:
  • Objects must contain at least $30$ LiDAR points.
  • Spatial positions: {front, behind, left, right, close by}
  • Relative size: {bigger than, smaller than}
  • Relative height: {taller than, shorter than}
Algorithm 1 Scene Graph Construction Pipeline
0: Scene graph information $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$
1: Initialize $\mathcal{G}\leftarrow\emptyset$.
2:for $D_{i}$ in $\mathcal{D}$ do
3:  Filter objects within sensor range and categories in $C$
4:  Further filter objects having more than 30 LiDAR points
5:  Obtain filtered boxes $B=\{b_{1},b_{2},\dots,b_{m}\}$
6:  Initialize relationship set $\mathcal{E}\leftarrow\emptyset$
7:  for each box pair $(b_{s},b_{o})$ in $B\times B$, where $s\neq o$ do
8:   Compute relationships $r\subseteq R$ using geometric attributes
9:   Update $\mathcal{E}\leftarrow\mathcal{E}\cup\{(s,r,o)\}$
10:  end for
11:  for each box $b_{s}$ in $B$ do
12:   Compute relationships $r_{ego}\subseteq R$ with ego node
13:   Update $\mathcal{E}\leftarrow\mathcal{E}\cup\{(s,r_{ego},ego)\}$
14:  end for
15:  Assemble scene graph $G_{i}\leftarrow(B,\mathcal{E})$
16:  Update $\mathcal{G}\leftarrow\mathcal{G}\cup\{G_{i}\}$
17:end for
18:return  $\mathcal{G}$
A.3 Training Configurations
4D Layout Generation.
The total loss combines both components:
$\boldsymbol{\delta}_{i}=\{(\Delta x_{i}^{\,t},\Delta y_{i}^{\,t})\}_{t=1}^{T}\,$,
where $T$ is the number of predicted future frames, set to 5 following the UniScene (Li et al. 2025) configuration. Each displacement is computed relative to the object’s position at the initial frame ($t=0$). A 1D U-Net serves as the denoising network, operating over the displacement sequence for each object node.
where $\mathbf{d}t$ is the noisy trajectory encoding at timestep $t$, $c$ denotes the conditioning input (i.e., node feature and bounding box), and $\varepsilon----{\theta}$ is the trajectory denoising network.
The total training loss is then given by:
(2) Denoising Architecture and Objective: The denoiser is a point-based U-Net following (Zheng et al. 2024b), which directly operates on normalized 3D point sets to capture both local and global structure efficiently.
Training uses the standard diffusion loss:
Range-based Point Cloud Generation.
$d_{i}^{\mathrm{norm}}=\frac{\log_{2}(d_{i}+1)}{\log_{2}(d_{\max}+1)}\,$, (17)
and then apply a linear rescaling to match the input range:
$x_{\mathrm{input}}=2\cdot x_{\mathrm{norm}}-1\,$, (18)
[…]
1 Joint $\mathbf{91.77}$ 87.41 10.57 35.43 $\mathbf{150}$
2 Split 90.72 $\mathbf{89.24}$ $\mathbf{9.79}$ $\mathbf{17.29}$ 270
@@ A.4 Evaluation Protocols
Scene-Level Evaluations.
To evaluate scene-level fidelity, we adopt standard distributional similarity metrics commonly used in LiDAR generative modeling (Zyrianov, Zhu, and Wang 2022). All metrics are computed over 10,000 generated samples for statistical reliability.
$\triangleright$ Fréchet Range Distance (FRD). To evaluate generation quality in the range-image domain, we compute the Fréchet Range Distance (FRD) by measuring the distance between the feature distributions of generated and real range-reflectance images. Following (Zyrianov, Zhu, and Wang 2022), features are extracted using RangeNet-53 (Milioto et al. 2019), a network pretrained for semantic segmentation on the nuScenes dataset (Caesar et al. 2020).
@@
$\triangleright$ Fréchet Point Distance (FPD). To assess generation quality in the 3D point cloud domain, we re-project generated range images back to point clouds and extract features using PointNet (Qi et al. 2017), pretrained for 16-class object classification on ShapeNet (Chang et al. 2015). Following (Shu, Park, and Kwon 2019), the Fréchet Point Distance (FPD) is computed as the distance between the feature distributions of generated and real point clouds:
@@
$\triangleright$ Jensen-Shannon Divergence (JSD). To evaluate 2D spatial distributions, we project point clouds into bird’s-eye view (BEV) occupancy maps following (Zyrianov, Zhu, and Wang 2022). For each sample, a 2D histogram over the BEV grid is computed. The Jensen–Shannon Divergence (JSD) is then calculated between the marginal distributions of generated and real BEV histograms:
@@
Object-Level Evaluations.
@@
Specifically, we apply a VoxelRCNN detector (Deng et al. 2021), pretrained on real nuScenes data, to the generated point clouds. We then compute standard 3D detection metrics (e.g., Average Precision, AP) by comparing the predicted boxes against the provided ground-truth bounding boxes used during conditioning. A higher CDA score indicates that the generated scenes are both geometrically accurate and semantically meaningful from the perspective of a pretrained downstream detector.
@@
A PointMLP classifier (Ma et al. 2022), pretrained on ground-truth foreground point clouds, is then used to classify each cropped region. The overall classification accuracy (OA) across all generated foregrounds is reported as the CFCA score. A higher CFCA indicates stronger semantic alignment with real-world object categories.
@@
$\triangleright$ Conditional Foreground Spatial Consistency (CFSC). CFSC evaluates the geometric fidelity of generated foreground point clouds under conditional generation. Inspired by GLENet (Zhang et al. 2023b) and illustrated in Fig. 8, it measures how well the generated points reconstruct the conditioning 3D bounding box.
@@
Temporal-Level Evaluations.
@@
Layout-Level Evaluations.
@@
Base Method Venue Representation Generator Architecture I.T. (s)
Uncond.
LiDARGen (Zyrianov, Zhu, and Wang 2022) ECCV’22 Range Image Energy-Based U-Net $67$
LiDM (Ran, Guizilini, and Wang 2024) CVPR’24 Range Image Latent DM Curve U-Net $3.5$
R2DM (Nakashima and Kurazume 2024) ICRA’24 Range Image DM Curve U-Net $3.5$
Cond.
UniScene (Li et al. 2025) CVPR’25 Voxel Volume Rendering Sparse U-Net $2.1$
OpenDWM (Ni et al. 2025b) CVPR’25 Voxel Latent DM U-Net $12.1$
OpenDWM-DiT (Ni et al. 2025b) CVPR’25 Voxel Latent DM Transformer $11.4$
LiDARCrafter Ours Range Image DM Curve U-Net $3.6$
A.5 Summary of Generations Baselines

### v3

@@ 3.3 Scene2Seq: Autoregressive LiDAR Synthesis
A core innovation of LiDARCrafter is its ability to generate the LiDAR stream autoregressively. In RGB video, textures and lighting change constantly, whereas a LiDAR sweep sees a mostly static environment, with only the ego vehicle and annotated objects moving. We exploit this stability by warping previously observed points to create a strong prior, as shown in Figure 4. Concretely, we back-project the first range image $\mathbf{I}^{0}$ to a point cloud $\mathbf{P}^{0}$, then split it with the layout boxes into background $\mathbf{B}^{0}$ and foreground sets $\{\mathbf{F}^{0}_{i}\}_{i=1}^{M}$. In later frames, we warp $\mathbf{B}^{0}$ with the ego pose, and update each $\mathbf{F}^{0}_{i}$ by its own motion prior, providing a strong drift-free geometric prior for the diffusion model at every denoising step.
@@
Dynamic-Object Warp. For each object $i$, we first shift its box center by its own cumulative planar offsets $(\Delta x_{i}^{t},\Delta y_{i}^{t})$, giving the world-frame position $\mathbf{u}_{i}^{t}=[x_{i}+\Delta x_{i}^{t},\;y_{i}+\Delta y_{i}^{t},\;z_{i}]^{\top}$. Its heading change $\psi_{i}^{t}$ is obtained exactly as in Equation 7. To express the box in the current ego frame, we apply the inverse ego transform: first translate by $\mathbf{u}_{0}^{t}$ and then rotate by $\psi_{0}^{t}$. The same rigid transform maps the stored foreground points $\mathbf{F}_{i}^{0}$ to $\mathbf{F}_{i}^{t}$. These warped foreground object points, combined with the updated background points, supply a strong geometric prior for the later timestep.
@@ 4 Experiments
4.1 Experimental Settings
We benchmark LiDARCrafter on the widely used nuScenes dataset [3], whose 32-beam LiDAR sweeps, 3D boxes, instance IDs, and HD maps provide rich conditioning cues. Evaluation combines classic static scores (FRD, FPD, JSD, MMD) with our object-, layout-, and motion-centric metrics from Section 3.4. Training uses six NVIDIA A40 GPUs, the three-branch layout diffuser runs $1M$ steps at batch $64$, and the range-image model ($32\times 1024$) runs $500K$ steps at batch $32$. All diffusers adopt $1024$ denoise steps for training and $256$ for sampling. More experimental settings appear in the appendix.
4.2 Scene-Level LiDAR Generation
Table 1: Evaluations of scene-level fidelity for LiDAR generation on the nuScenes dataset. MMD values are reported in $10^{-4}$ and JSD in $10^{-2}$. Lower is better for all metrics.
@@ 4.2 Scene-Level LiDAR Generation
Table 5: Comparison of object generation consistency using CFCA ($\uparrow$) and CFSC ($\uparrow$). CFCA measures classification accuracy on generated points using a PointMLP trained on real data. CFSC assesses geometric consistency by regressing boxes from generated points and computing IoU. Numbers indicate the point count within each box.
@@
Table 6: Comparison of temporal consistency in 4D LiDAR generation using TTCE ($\downarrow$) and CTC ($\downarrow$). TTCE measures the transformation error from the point cloud registration perspective, indicating the consistency of temporal changes; CTC computes the Chamfer Distance between the adjacent generated LiDAR frames. Numbers indicate the frame intervals.
@@
Table 7: Ablation on foreground conditioning methods for generation. 2D masks are projected from 3D boxes.
@@
Whole-Scene Fidelity. Table 2 compares LiDARCrafter against recent LiDAR generation baselines in terms of whole-scene fidelity. Our method outperforms all existing methods, achieving the lowest FRD ($194.37$, a $20\%$improvement over R2DM) and FPD ($8.64$), indicating more accurate reconstruction of both 2D range structures and 3D point distributions. As shown in Figure 5, our method produces scans that closely resemble the ground truth, with well-preserved foreground structures, while other methods suffer from background noise or blurring.
Foreground Object Accuracy. To evaluate the fidelity of synthesized foregrounds, we apply a pre-trained VoxelRCNN detector [5] to the generated scenes. We report Foreground Detection Confidence (FDC), which reflects the detector’s confidence (cf. Table 2), and Conditioned Detection Accuracy (CDA), which measures the average precision (AP) of detected boxes (cf. Table 4). LiDARCrafter achieves the highest FDC scores across most categories, indicating superior object fidelity, and also attains the best detection AP, demonstrating stronger alignment between synthesized structures and the given conditions.
@@ 4.3 Object-Level LiDAR Generation
Semantic and Geometric Consistency. To further evaluate object quality under conditioning, we introduce two metrics: CFCA for semantic fidelity and CFSC for geometric consistency (Table 6). For semantic fidelity, we apply a PointMLP [27] classifier trained on real data to classify generated instances, yielding a CFCA score of $73.48\%$ for LiDARCrafter, indicating strong alignment with real-world categories. For geometric consistency, we use a conditional variational autoencoder to regress bounding boxes from generated point clouds, and compute the mean IoU with ground truth. LiDARCrafter achieves the highest IoU across all point count settings, demonstrating superior adherence to geometric constraints.
@@ 4.4 Autoregressive 4D LiDAR Generation
Temporal Consistency. We evaluate temporal consistency in 4D LiDAR generation in Table 6. TTCE measures the error between the predicted and ground-truth transformation matrices obtained via point cloud registration, while CTC computes the Chamfer Distance between consecutive frames. Our approach achieves the lowest TTCE scores across both frame intervals and maintains competitive CTC performance at all intervals, demonstrating strong temporal coherence. Qualitative comparisons in Figure 6 further show that LiDARCrafter produces sequences with consistent structure and fine geometric detail, whereas other methods often suffer from degraded fidelity over time.
@@ 4.6 Applications
Leveraging its object-centric generation capability, LiDARCrafter enables the synthesis of rare and diverse corner-case scenarios, which are valuable for data augmentation and for evaluating the robustness of downstream algorithms. As illustrated in Figure 7, our method can generate challenging situations such as a vehicle cutting into the driving lane, a pedestrian emerging in a bus blind spot, a collision between two adjacent cars, and an overtaking maneuver. These results demonstrate that our method not only produces the desired scenarios on demand, but also maintains strong temporal coherence throughout the sequence. Additional qualitative examples are provided in the appendix.
@@ 5 Conclusion
6 Additional Implementation Details
@@
  • Notation Summary (Section 6.1): We summarize the notations used throughout the main manuscript for clarity.
  • Scene Graph Construction (Section 6.2): We explain how scene graphs are built from dataset annotations, which serve as training data for LiDARCrafter.
  • Training Pipeline (Section 6.3): We detail the data preprocessing steps, loss functions, and hyperparameter settings used for both 4D layout and point cloud generation.
  • Evaluation Suite (Section 6.4): We clarify the motivation behind each proposed metric and specify the exact computation procedures.
  • Baseline Models (Section 6.5): We briefly describe the baseline methods used for comparisons.
6.1 Summary of Notations
For better readability, the notations used in this work have been summarized in Table 11.
6.2 Scene Graph Construction from Annotations
To train LiDARCrafter, we construct an ego-centric scene graph for each LiDAR frame using annotations from the nuScenes dataset. Each scene graph consists of nodes representing the ego vehicle and nearby foreground objects, along with directed edges that capture their spatial and semantic relationships. The construction pipeline is summarized in Algorithm 1, and detailed as follows:
  • Objects must contain at least $30$ LiDAR points.
  • Spatial positions: {front, behind, left, right, close by}
  • Relative size: {bigger than, smaller than}
  • Relative height: {taller than, shorter than}
Algorithm 1 Scene Graph Construction Pipeline
0: Scene graph information $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$
1: Initialize $\mathcal{G}\leftarrow\emptyset$.
2:for $D_{i}$ in $\mathcal{D}$ do
3:  Filter objects within sensor range and categories in $C$
4:  Further filter objects having more than 30 LiDAR points
5:  Obtain filtered boxes $B=\{b_{1},b_{2},\dots,b_{m}\}$
6:  Initialize relationship set $\mathcal{E}\leftarrow\emptyset$
7:  for each box pair $(b_{s},b_{o})$ in $B\times B$, where $s\neq o$ do
8:   Compute relationships $r\subseteq R$ using geometric attributes
9:   Update $\mathcal{E}\leftarrow\mathcal{E}\cup\{(s,r,o)\}$
10:  end for
11:  for each box $b_{s}$ in $B$ do
12:   Compute relationships $r_{ego}\subseteq R$ with ego node
13:   Update $\mathcal{E}\leftarrow\mathcal{E}\cup\{(s,r_{ego},ego)\}$
14:  end for
15:  Assemble scene graph $G_{i}\leftarrow(B,\mathcal{E})$
16:  Update $\mathcal{G}\leftarrow\mathcal{G}\cup\{G_{i}\}$
17:end for
18:return $\mathcal{G}$
6.3 Training Configurations
6.3.1 4D Layout Generation.
The total loss combines both components:
$\boldsymbol{\delta}_{i}=\{(\Delta x_{i}^{\,t},\Delta y_{i}^{\,t})\}_{t=1}^{T}\,$,
where $T$ is the number of predicted future frames, set to 5 following the UniScene [19] configuration. Each displacement is computed relative to the object’s position at the initial frame ($t=0$). A 1D U-Net serves as the denoising network, operating over the displacement sequence for each object node.
where $\mathbf{d}t$ is the noisy trajectory encoding at timestep $t$, $c$ denotes the conditioning input (i.e., node feature and bounding box), and $\varepsilon_{\theta}$ is the trajectory denoising network.
The total training loss is then given by:
(2) Denoising Architecture and Objective: The denoiser is a point-based U-Net following [65], which directly operates on normalized 3D point sets to capture both local and global structure efficiently.
Training uses the standard diffusion loss:
Base Method Venue Representation Generator Architecture I.T. (s)
Uncond.
LiDARGen [69] ECCV’22 Range Image Energy-Based U-Net $67$
LiDM [37] CVPR’24 Range Image Latent DM Curve U-Net $3.5$
R2DM [30] ICRA’24 Range Image DM Curve U-Net $3.5$
Cond.
UniScene [19] CVPR’25 Voxel Volume Rendering Sparse U-Net $2.1$
OpenDWM [33] CVPR’25 Voxel Latent DM U-Net $12.1$
OpenDWM-DiT [33] CVPR’25 Voxel Latent DM Transformer $11.4$
LiDARCrafter Ours Range Image DM Curve U-Net $3.6$
[…]
1 Joint $\mathbf{91.77}$ 87.41 10.57 35.43 $\mathbf{150}$
2 Split 90.72 $\mathbf{89.24}$ $\mathbf{9.79}$ $\mathbf{17.29}$ 270
6.3.2 Range-based Point Cloud Generation.
$d_{i}^{\mathrm{norm}}=\frac{\log_{2}(d_{i}+1)}{\log_{2}(d_{\max}+1)}\,$, (17)
and then apply a linear rescaling to match the input range:
$x_{\mathrm{input}}=2\cdot x_{\mathrm{norm}}-1\,$, (18)
@@ 6.4 Evaluation Protocols
6.4.1 Scene-Level Evaluations.
To evaluate scene-level fidelity, we adopt standard distributional similarity metrics commonly used in LiDAR generative modeling [69]. All metrics are computed over 10,000 generated samples for statistical reliability.
$\triangleright$ Fréchet Range Distance (FRD). To evaluate generation quality in the range-image domain, we compute the Fréchet Range Distance (FRD) by measuring the distance between the feature distributions of generated and real range-reflectance images. Following [69], features are extracted using RangeNet-53 [29], a network pretrained for semantic segmentation on the nuScenes dataset [3].
@@
$\triangleright$ Fréchet Point Distance (FPD). To assess generation quality in the 3D point cloud domain, we re-project generated range images back to point clouds and extract features using PointNet [35], pretrained for 16-class object classification on ShapeNet [4]. Following [40], the Fréchet Point Distance (FPD) is computed as the distance between the feature distributions of generated and real point clouds:
@@
$\triangleright$ Jensen-Shannon Divergence (JSD). To evaluate 2D spatial distributions, we project point clouds into bird’s-eye view (BEV) occupancy maps following [69]. For each sample, a 2D histogram over the BEV grid is computed. The Jensen–Shannon Divergence (JSD) is then calculated between the marginal distributions of generated and real BEV histograms:
@@
6.4.2 Object-Level Evaluations.
@@
Specifically, we apply a VoxelRCNN detector [5], pretrained on real nuScenes data, to the generated point clouds. We then compute standard 3D detection metrics (e.g., Average Precision, AP) by comparing the predicted boxes against the provided ground-truth bounding boxes used during conditioning. A higher CDA score indicates that the generated scenes are both geometrically accurate and semantically meaningful from the perspective of a pretrained downstream detector.
@@
A PointMLP classifier [27], pretrained on ground-truth foreground point clouds, is then used to classify each cropped region. The overall classification accuracy (OA) across all generated foregrounds is reported as the CFCA score. A higher CFCA indicates stronger semantic alignment with real-world object categories.
@@
$\triangleright$ Conditional Foreground Spatial Consistency (CFSC). CFSC evaluates the geometric fidelity of generated foreground point clouds under conditional generation. Inspired by GLENet [62] and illustrated in Fig. 8, it measures how well the generated points reconstruct the conditioning 3D bounding box.
@@
6.4.3 Temporal-Level Evaluations.
@@
6.4.4 Layout-Level Evaluations.
@@
6.5 Summary of Generations Baselines